OpenStudy (anonymous):

Solve for x: 5 over quantity x squared minus 4 plus 2 over x equals 2 over quantity x minus 2 x = 8 x = –4 x = 8 and x = –4 No Solution i will reward medal

OpenStudy (anonymous):

the equation looks like \[\frac{ 5 }{ x ^{2}-4}+\frac{ 2 }{ x }=\frac{ 2 }{ x-2 }\]

OpenStudy (anonymous):

whats the first step? did you solve that last one?

OpenStudy (anonymous):

i got the last one solved

OpenStudy (anonymous):

and again you find the common denominator but i get confused after that

OpenStudy (anonymous):

@dalia_lam pleaze help. i will give you another medal

OpenStudy (anonymous):

What is the common denominator? Can you see it?

OpenStudy (anonymous):

it is x-4.after i get the denominator is when i get confused

OpenStudy (anonymous):

the answer is in blue ink. sorry for the messy paper :P

OpenStudy (anonymous):

the answer is 8